OpenStudy (anonymous):

4-4/x+1 /4+ 4/x simply complex fraction assume no division by 0.

OpenStudy (anonymous):

terms -4/x and 4/x got cancelled and u left with 4 + 1/4 = 17/4 is ur answer

OpenStudy (anonymous):

please explain this

OpenStudy (anonymous):

see : 4-4/x + 1/4 +4/x ( u need to work on the constant terms and terms with variables) so here these : 4 + 1/4 ( constant terms ) and -4/x + 4/x ( are terms with variable x ) as variable terms are of opposite signs viz. one is positive and another one is negative so they got cancelled and u can add other constant terms which gives u : (16+1)/4 = 17/4